B2H6 + 3O2 -> 2Hbo2 + 2H2O

What mass of water us produced if 1.38 moles of HBO2 are produced ?

How many moles of water are produced from 4.75 moles of oxygen?

How many moles of B2H6 are needed to react to produce 7.90 grams of water?
1
Expert's answer
2020-12-15T07:52:38-0500

What mass of water us produced if 1.38 moles of HBO2 are produced ?

n(H2O) = n(HBO2) = 1.38 mol

m(H2O) = "1.38\;mol \times 18 \;g/mol = 24.84 \;g"

How many moles of water are produced from 4.75 moles of oxygen?

n(H2O) "= \frac{2}{3}n(O_2) = \frac{2}{3} \times 4.75 = 3.16 \;mol"

How many moles of B2H6 are needed to react to produce 7.90 grams of water?

n(H2O) ="\frac{7.90}{18.0} = 0.438 \;mol"

n(B2H6) "= \frac{1}{2}n(H_2O) = \frac{1}{2} \times 0.438 = 0.219 \;mol"